OpenStudy (anonymous):

Simplify the rational expression: x^2 - y^2/(x + y)

OpenStudy (aum):

Is it \(\Large \frac{x^2-y^2}{x+y} \) or \(\Large x^2 - \frac{y^2}{x+y} \) ?

OpenStudy (anonymous):

\[\frac{ x ^{2} - y ^{2}}{ (x + y) }\]